OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

is it to find x and y??

OpenStudy (anonymous):

psh got me it has these multiple questions (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

the answer is x=12 and y=-1 just use substitution method

OpenStudy (anonymous):

can you help me on this one too pwese ? (: Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1

OpenStudy (anonymous):

2x+y=1---1 4x+2y=-1 ----2 2x+y=1 y=1-2x-----3 sub 3 into 2 4x+2y=-1 4x+2(1-2x)=-1 solve it and u get x